A firm’s marginal cost function is Q2 + 2Q+ 4. Find the total cost function if the fixed costs are 100
Answer #1

Answer

Formula:

where TC = Total Cost and MC = Marginal Cost

=> TC = Q3/3 + 2Q2/2 + 4Q + F

=> TC = Q3/3 + Q2 + 4Q + F

where F = is some constant and is a Fixed cost and here Fixed cost = 100

=> Total Cost(TC) = Q3/3 + Q2 + 4Q + 100